### v1

# Train Once for All: A Transitional Approach for Efficient Aspect Sentiment Triplet Extraction

Abstract
Aspect-Opinion Pair Extraction (AOPE) and Aspect Sentiment Triplet Extraction (ASTE) have gained significant attention in natural language processing. However, most existing methods are a pipelined framework, which extracts aspects/opinions and identifies their relations separately, leading to a drawback of error propagation and high time complexity. Towards this problem, we propose a transition-based pipeline to mitigate token-level bias and capture position-aware aspect-opinion relations. With the use of a fused dataset and contrastive learning optimization, our model learns robust action patterns and can optimize separate subtasks jointly, often with linear-time complexity. The results show that our model achieves the best performance on both the ASTE and AOPE tasks, outperforming the state-of-the-art methods by at least 6.98% in the F1 measure. The code is available at https://github.com/Paparare/trans_aste.
Train Once for All: A Transitional Approach for Efficient Aspect Sentiment Triplet Extraction
Xinmeng Hou1  Lingyue Fu2  Chenhao Meng2  Hai Hu2thanks: Corresponding author 1Columbia University, New York, NY, USA 2Shanghai Jiao Tong University, Shanghai, China fh2450@tc.columbia.edu fulingyue@sjtu.edu.cn, chenhaomeng@sjtu.edu.cn, hu.hai@sjtu.edu.cn
1 Introduction
Aspect-Based Sentiment Analysis (ABSA) is a fine-grained sentiment analysis task that identifies specific aspects in text and analyzes the sentiments linked to them. It involves subtasks such as: Aspect Extraction (AE): Identifying mentioned aspects. Opinion Extraction (OE): Extracting opinion terms related to aspects. Aspect-Opinion Pair Extraction (AOPE): Extracting aspect-opinion pairs. Aspect-Sentiment Triplet Extraction (ASTE): Identifying aspect, opinion, and sentiment polarity triplets.
Figure 1: For the sentence ’Gourmet food is delicious. Good service, but not so welcoming,’ this example demonstrates the processing steps in both classic and transitional methods for extracting aspect-opinion pairs and tagging sentiment polarity. AE stands for Aspect Extraction, OE for Opinion Extraction, POS represents positive sentiment, and NEG represents negative sentiment.
Classic ABSA approaches often use a pipeline architecture, performing subtasks like aspect extraction (AE) and opinion extraction (OE) separately. For instance, in the sentence: “Gourmet food is delicious. Good service, but not so welcoming.,” AE identifies gourmet food and service as aspects, while OE extracts delicious, good, and not so welcoming as opinions. These outputs are combined to form aspect-opinion pairs, with sentiment tagging assigning polarities to create triplets. Past studies have generally achieved satisfactory performance on aspect and opinion extraction Jiang et al. (2023); Wang et al. (2023); Chakraborty (2024). Adding the according sentiment tags with pair, ASTE is the most integrated task for aspect-based sentiment analysis that has seen the development of diverse models leveraging various methodologies. Pipeline-based approaches, such as Peng-Two-stage Peng et al. (2020), decompose the task into multiple stages for modular refinement. Sequence-to-sequence frameworks like BARTABSA Yan et al. (2021) employ pretrained transformers to generate triplets flexibly. Sequence-tagging methods, including GTS Wu et al. (2020) and JET-BERT Xu et al. (2020), annotate tokens for precise identification of relationships. Machine Reading Comprehension (MRC)-based models, such as COM-MRC Zhai et al. (2022) and Triple-MRC Zou et al. (2024), reframe the task as query answering for efficient extraction. Graph-based approaches such as EMC-GCN Chen et al. (2022), BDTF Chen et al. (2022), and DGCNAP Li et al. (2023) use graph structures to capture semantic and syntactic interactions. Tagging schema-based models, exemplified by STAGE-3D Liang et al. (2023), use hierarchical schemas for multi-level extraction, while lightweight models like MiniConGTS Sun et al. (2024) focus on efficiency with reduced computational costs.
Although previous studies have shown satisfactory results for individual entity extraction, performance on pair and triplet extraction remains suboptimal, with state-of-the-art (SOTA) scores in the 60s and 70s Sun et al. (2024). Two key challenges contribute to this gap: (1) Disconnected Aspect-Opinion Extraction: Opinions are often extracted independently from their corresponding aspects in previous studies Liang et al. (2023); Sun et al. (2024). While relationships can be added as an auxiliary factor to assist pair extraction Liu et al. (2022); Wang et al. (2023), this approach loses critical contextual information by treating aspects and opinions as separate entities. This limits the effectiveness of many token-based extraction methods. (2) Time Complexity with Longer Sequences: Methods using 2D matrix tagging Liang et al. (2023); Sun et al. (2024) to capture relationships between tokens face significant increases in time complexity as the length of the token sequence increases. This computational burden restricts their scalability, especially for longer texts in practical applications.
As an overview, the critical contributions of our study are as follows:
Efficient Transition Pipeline for AOPE:
We propose a transition pipeline for simultaneous aspect-opinion pair extraction, achieving linear time complexity $O(n)$.
Augmented Learning for Model Optimization:
Building on this transitional backbone, we introduce a contrastive-augmented optimization method to enhance model efficiency.
New SOTA Performance on AOPE and ASTE:
Our model achieves new state-of-the-art (SOTA) performance on both AOPE and ASTE tasks.
2 Related Work
2.1 Transition-based Methods
Transition-based approaches are widely used in dependency parsing, leveraging shift-reduce and bidirectional arc actions (left-arc, right-arc) for efficient $O(n)$ parsing Aho and Ullman (1973); Nivre (2003); Cer et al. (2010). These parsers maintain stack, buffer, and arc relations to track transitions and have been extended to tasks like token segmentation Zhang et al. (2016) and argument mining Bao et al. (2021). Beyond parsing, transition-based methods have been applied to structured sentiment analysis, generating graph structures for opinion extraction Fernández-González (2023) and incremental constituency parsing using graph neural networks Yang and Deng (2020). A unifying theory bridges transition-based parsing and sequence labeling for simplified learning Gómez-Rodríguez et al. (2020), while in AMR parsing, structure-aware fine-tuning achieves state-of-the-art results Zhou et al. (2021). Their flexibility and efficiency make them effective across diverse NLP tasks.
2.2 Contrastive-based Optimization
Contrastive learning has demonstrated its effectiveness in various domains, achieving state-of-the-art results in token-independent extraction tasks with models such as MINIconGTS Sun et al. (2024). Recent works have further explored its potential: contrastive learning for prompt-based few-shot language learners enhances generalization through augmented "views" Jian et al. (2022); contrastive learning as goal-conditioned reinforcement learning provides a novel loss function with theoretical guarantees for improved success in goal-directed tasks Eysenbach et al. (2023), and studies analyzing the role of margins reveal critical factors like positive sample emphasis for better generalization Rho et al. (2023). These advances showcase contrastive learning’s flexibility and robustness, which inspired us to integrate it into our transition-based AOPE and ASTE to refine representation learning and improve model performance.
3 Methodology
Action Symbolic Expression
Shift ($SF$) $(\sigma_{0},\,\beta_{0}\mid\beta_{1},E,C,R)\xrightarrow{SH}(\sigma_{0}\mid\sigma_{1},\beta_{1},E,C,R)$
Stop ($ST$) $(\sigma_{0},\O,A,O,R)\xrightarrow{ST}(\O,\O,A,O,R)$
[…]
Left Constituent Removal ($L_{n}$) $(\sigma_{0}\mid\sigma_{1},\,\beta_{0},A,O,R)\xrightarrow{L_{n}}(\sigma_{1},\,\beta_{0},A,O,R)$
Right Constituent Removal ($R_{n}$) $(\sigma_{0}\mid\sigma_{1},\,\beta_{0},A,O,R)\xrightarrow{R_{n}}(\sigma_{1},\,\beta_{0},A,O,R)$
Left-Relation Formation ($LR$) $(\sigma_{0}\mid\sigma_{1},\beta_{0},A,O,R)\xrightarrow{LR}(\sigma_{0}\mid\sigma_{1},\beta_{0},A\cup\sigma_{1},O\cup\sigma_{0},R\cup\sigma_{0}\xleftarrow{}\sigma_{1})$
Right-Relation Formation ($RR$) $(\sigma_{0}\mid\sigma_{1},\beta_{0},E,C,R)\xrightarrow{RR}(\sigma_{0}\mid\sigma_{1},\beta_{0},A\cup\sigma_{0},O\cup\sigma_{0},R\cup\sigma_{0}\xrightarrow{}\sigma_{1})$
Table 1: Symbolic Expressions for the Proposed Actions.
In this section, we introduce a novel paradigm that integrates the extraction of aspect-opinion pairs into a process reminiscent of constructing a parsing-directed graph. This model progressively constructs and annotates aspect-oriented relations using input sequences enriched with contextual features. Unlike the classic approach, which employs three sets to record action histories, we utilize five sets, including stack, buffer, aspect, and pair, to show how aspect and opinion are extracted sitimulately.
3.1 Transitional Operations and State Change
Phrase Action Stack ($\sigma$) Buffer ($\beta$) Aspect Opinion Pair
[] [$\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$]
1 $SF$ [$\sigma_{1}$] [$\beta_{2}$, $\beta_{3}$, $\beta_{4}$]
2 $SF$ [$\sigma_{1}$, $\sigma_{2}$] [$\beta_{3}$, $\beta_{4}$]
3 $M$ [$\sigma_{1\&2}$] [$\beta_{3}$, $\beta_{4}$]
4 $SF$ [$\sigma_{1\&2}$, $\sigma_{3}$] [$\beta_{4}$]
5 $R_{n}$ [$\sigma_{1\&2}$] [$\beta_{4}$]
6 $SF$ [$\sigma_{1\&2}$, $\sigma_{4}$] []
7 $RR$ [$\sigma_{1\&2}$, $\sigma_{4}$] [] [$\sigma_{1\&2}$] [$\sigma_{4}$] ($\sigma_{1\&2}\rightarrow\sigma_{4}$)
9 $ST$ [] [] [$\sigma_{1\&2}$] [$\sigma_{4}$] ($\sigma_{1\&2}\rightarrow\sigma_{4}$)
Table 2: State changes for "Gourmet food is delicious" using symbolic representation. Here, $\sigma_{1}$ corresponds to "Gourmet", $\sigma_{2}$ to "food", $\sigma_{3}$ to "is", and $\sigma_{4}$ to "delicious". Similarly, $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, and $\beta_{4}$ correspond to tokens in the buffer in sequence.
The construction of relationships between phrases is achieved through systematic transition actions, conceptualized as directed edges linking two nodes (tokens), $N_{1}$ and $N_{2}$. A relationship transitioning $N_{1}$ to $N_{2}$ under relation $l_{1}$ is denoted as $RR=N_{1}\xrightarrow{l_{1}}N_{2}$ or $LR=N_{1}\xleftarrow{l_{1}}N_{2}$, representing aspect-oriented direction. To interpret causal links, $l_{1}$ is defined as $l_{L}$ or $l_{R}$ for bidirectional relations. Opinion constituents ($E_{1}$) may precede or follow aspect constituents, which can span multiple tokens (e.g., “gourmet food,” “not bad”). This requires additional steps to consolidate multi-token constituents. For the ASTE task, we define seven distinct transition actions that combine token retrieval, termination, and token merging.
The action framework aims to prevent information leakage and ensure flexibility in relation arching. Actions are recorded as tuples $T=(\sigma,\beta,A,O,R)$, representing the stack, buffer, aspect, opinion, and relation. The ASTE task includes default actions, which are universal and required for all data, primary actions that handle merging or removal, and secondary actions for relation arching, which depend on the outcomes of the primary actions. Verbal and symbolic representations of state transitions for each action are provided below.
Default Actions:
  1. 1.
    Shift ($SF$) moves a token from the tokenized stack into the buffer for further processing.
  2. 2.
    Stop ($ST$) halts the process when only one token remains in the buffer, and the stack is empty.
Primary Actions:
  1. 1.
    Merge ($M$) combines multiple tokens in the buffer into a single compound target.
  2. 2.
    Left Constituent Removal ($L_{n}$) removes the left constituent from the buffer.
  3. 3.
    Right Constituent Removal ($R_{n}$) removes the right constituent from the buffer.
Secondary Actions:
  1. 1.
    Left-Relation Formation ($LR$) creates a relation from the right aspect constituent to the left opinion constituent.
  2. 2.
    Right-Relation Formation ($RR$) creates a relation from the left aspect constituent to the right opinion constituent.
Table 1 provides a symbolic illustration of how the symbolic state is constructed and utilized. Take the sentence "Gourmet food is delicious" as an example. Table 2 demonstrates the process of moving tokens from the buffer to the stack, deciding whether they should be merged into a single entity or removed, and finally evaluating them for relation formation. It is important to note that the set of actions shown in the figure is not the only way to extract the "Gourmet food" and "delicious" aspect-opinion pair. The alternative solution for parsing the sentence is shown in Appendix 5.
3.2 Trans-AOPE State Representation
Figure 2: The complete process of the transition-based model is illustrated. Purple highlights represent the transition-based pair extraction, while orange indicates the final step of sentiment tagging.
The model we propose consists of two core stages: pair extraction with a designed transitional action slot (in purple) and pair-based sentiment tagging (in orange), as illustrated in Figure 2.
In the first stage, the input, denoted as $I^{n}_{1}=(t_{1},t_{2},\ldots,t_{n})$, is a sequence of tokens. The output is a sequence of actions, represented as $A^{m}_{1}=(a_{1},a_{2},\ldots,a_{m})$. This process can be conceptualized as a search for the optimal action sequence, $A^{*}$, given the input sequence $I^{n}_{1}$. At each step $n$, the model predicts the next action based on the current system state, $S$, and the sequence of prior actions, $A^{n-1}_{1}$ The updated system state, $S_{n+1}$ , is determined by the specific action $a_{t}$. We define $r_{n}$ as a symbolic representation for calculating the probability of the action $a_{n}$ at step $n$. This probability is computed as follows:
$p(a_{n}|r_{n})=\frac{\exp(w_{a_{n}}^{\top}r_{n}+b_{a_{n}})}{\sum\nolimits_{a^{\prime}\in\mathcal{A}(S)}{\exp(w_{a^{\prime}}^{\top}r_{n}+b_{a^{\prime}})}}$ (1)
Here, $w_{a}$ is a learnable parameter vector, and $b_{a}$ is a bias term. The set $\mathcal{A}(S)$ represents the legal actions available given the current parser state. The overall optimization objective for the model is defined as:
$(A^{*},S^{*})$ $=\underset{{}_{A,S}}{\mathrm{argmax}}\prod_{n}p(a_{n},S_{n+1}|A^{n-1}_{1},S_{n})$ (2)
$=\underset{{}_{A,S}}{\mathrm{argmax}}\prod_{n}p(a_{n}|r_{n})$
In this stage, the ASTE task is incorporated into a transition-based action prediction task. In order to decode efficiently, the action boasting the maximum probability is chosen through a greedy approach until the parsing procedure reaches a point of termination. This model thus forms the basis of an efficient transition-based prediction system that utilizes a representation of the current system state and action history to generate the most likely subsequent action at each step in a sequence. By integrating these concepts, the model successfully carries out parsing tasks while avoiding information leakage and ensuring flexibility in the relation arching process.
3.3 Transition Implementation with Neural Model
This section outlines the transition-based parsing process using neural models. RoBERTa Liu et al. (2019) encodes textual phrases, while UniLSTM Hochreiter and Schmidhuber (1997) and BiLSTM Graves and Schmidhuber (2005) model dynamic transitions. The parser state is initialized and evolves through a sequence of actions with distributed embeddings and LSTMs representing the state at each step. At the end stage, a Multi-Layer Perceptron (MLP) is employed to classify the sentiment associated with each pair or triplet based on the evolved parser state.
Token representations
Consider the process of parsing a text $d^{n}_{1}$ = ($p_{1}$, $p_{2}$, $\ldots$, $p_{n}$) , consisting of $n$ phrases. Each phrase $p_{i}$ = ($w_{i1}$, $w_{i2}$, $\ldots$, $w_{il}$) contains $l$ tokens. A phrase can be represented as a sequence $x_{i}$ = ($\mathrm{[CLS]}$, $t_{i1}$, $\ldots$, $t_{il}$, $\mathrm{[SEP]}$), where $\mathrm{[CLS]}$ is a special classification token whose final hidden state serves as the aggregate sequence feature, and $\mathrm{[SEP]}$ is a separator token. The hidden representation of each phrase is computed as $h_{p_{i}}$ = $\mathrm{RoBERTa}$($x_{i}$) $\in\mathbb{R}^{d_{b}\times|l_{i}|}$, where $d_{b}$ is the hidden dimension size, and $|l_{i}|$ is the length of the sequence $x_{i}$. Finally, the entire text $d^{n}_{1}$ is represented as a list of tokens: $h_{d}$ = [$h_{p_{1}}$, $h_{p_{2}}$, $\ldots$, $h_{p_{n}}$].
State Initialization
At the start of the parsing process, the parser’s state is initialized as $(\beta=\varnothing,\sigma=[1,2,\ldots,n],E=\varnothing,C=\varnothing,R=\varnothing)$, where $\sigma$ is the stack, $\beta$ is the buffer, and $E$, $C$, and $R$ are empty sets representing different outputs. The state evolves through a sequence of actions, progressively consuming elements from the buffer $\beta$ and constructing the output. This process continues until the parser reaches its terminal state when there is only one token left in buffer, represented as $(\beta=[SEP],\sigma=\varnothing,E,C,R)$.
Step-by-Step Parser State Representation.
For the action sequence, each action $a$ is mapped to a distributed representation $e_{a}$ through a lookup table $E_{a}$. An unidirectional LSTM is then utilized to capture the complete history of actions in a left-to-right manner at each step $t$:
$\alpha_{t}=\text{LSTM}_{a}(a_{0},a_{1},\ldots,a_{t-1},a_{t})$ (3)
Upon generation of a new action $a_{t}$, its corresponding embedding $e_{a_{t}}$ is integrated into the rightmost position of LSTMa. To further refine the representation of the pair $(\sigma_{1},\sigma_{0})$, their relative positional distance $d$ is also encoded as an embedding $e_{d}$ from a lookup table $E_{d}$. The composite representation of the parser state at step $t$ encompasses these varied features.
The parser state is represented as a triple $(\beta_{s},\sigma_{s},A_{t})$, where $\sigma_{s}$ denotes the stack sequence $(\sigma_{0},\sigma_{1},\ldots,\sigma_{n})$, $\beta_{s}$ represents the buffer sequence $(\beta_{0},\beta_{1},\ldots,\beta_{n})$, and $A_{t}$ encapsulates the action history $(a_{0},a_{1},\ldots,a_{t-1},a_{t})$. The stack ($\sigma_{n}$) and buffer ($\beta_{n}$) are encoded using bidirectional LSTMs as follows:
$=\text{BiLSTM}((\sigma_{n},\beta_{0}),(\sigma_{n-1},\beta_{1}),$ (4)
$\ldots,(\sigma_{0},\beta_{n}))$
Here, $s_{t}$ and $b_{t}$ are the output feature representations of the stack and buffer, respectively. Each of these representations consists of forward and backward components: $\sigma_{t}=(\overrightarrow{\sigma_{t}},\overleftarrow{\sigma_{t}})$ and $\beta_{t}=(\overrightarrow{\beta_{t}},\overleftarrow{\beta_{t}})$. The forward and backward components are matrices in $\mathbb{R}^{d_{l}\times|\sigma_{t}|}$ and $\mathbb{R}^{d_{l}\times|\beta_{t}|}$, respectively, where $d_{l}$ is the hidden dimension size of the LSTM, and $|\sigma_{t}|$, $|\beta_{t}|$ are the lengths of the sequences $\sigma_{t}$ and $\beta_{t}$.
3.4 Optimization Implementation
We compare two optimization strategies: regular optimization using Cross-Entropy Loss and contrastive-based optimization, which aligns predicted and true action embeddings. A weight study will evaluate the impact of the positioning of two components in augmented optimization on model performance.
3.4.1 Base Optimization
Both action and sentiment classification tasks are optimized using the Cross-Entropy Loss as base optimization, defined as follows:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}y_{i}^{j}\log(p_{i}^{j}),$ (5)
where $N$ is the number of samples, $M$ is the number of classes (either action or sentiment classes), $y_{i}^{j}$ is a binary indicator (0 or 1) indicating whether class $j$ is the correct class for sample $i$, and $p_{i}^{j}$ is the predicted probability for class $j$ for sample $i$. The total loss is the sum of the losses for both tasks:
$\mathcal{L}_{\text{base}}=\mathcal{L}_{\text{action}}+\mathcal{L}_{\text{sentiment}}.$ (6)
3.4.2 Contrastive-based Optimization
Given the predicted action logits $\mathbf{A}_{\text{logits}}$ and the true action labels $\mathbf{A}_{\text{true}}$, the predicted actions are obtained by applying a softmax function followed by an argmax:
$\mathbf{A}_{\text{pred}}=\text{argmax}(\text{softmax}(\mathbf{A}_{\text{logits}}))$
The predicted and true actions are then embedded as follows:
$\mathbf{E}_{\text{pred}}=\text{Embed}(\mathbf{A}_{\text{pred}}),\quad\mathbf{E}_{\text{true}}=\text{Embed}(\mathbf{A}_{\text{true}})$
For each sample $i$, compute the cosine similarity between $\mathbf{E}_{\text{pred}}^{(i)}$ and all $\mathbf{E}_{\text{true}}^{(j)}$ to form a similarity matrix. Using this matrix, the diagonal elements represent positive pairs, while the non-diagonal elements represent negative pairs.
Next, calculate the positive and negative cosine similarities, and apply the exponential function:
$e_{\text{pos}}=\exp(\text{cos\_sim}(\mathbf{E}_{\text{pred}},\mathbf{E}_{\text{true}})\cdot\mathbf{M}_{\text{pos}})$ (7)
$e_{\text{all}}=\exp(\text{cos\_sim}(\mathbf{E}_{\text{pred}},\mathbf{E}_{\text{true}}))$ (8)
where, cos_sim denote cosine similarity function $\mathbf{M}_{\text{pos}}$ is a mask that selects only positive pairs (diagonal entries).
Now compute the contrastive loss as:
$\mathcal{L}_{\text{con}}=-\frac{1}{N}\sum_{j=1}^{N}\log\left(\frac{\exp_{\text{pos}}^{(j)}}{\exp_{\text{all}}^{(j)}}\right)$ (9)
Finally, the total loss is computed as:
$\mathcal{L}_{\text{total}}=\omega_{1}\cdot\mathcal{L}_{\text{base}}+\omega_{2}\cdot\mathcal{L}_{\text{con}}$ (10)
where $\omega$ are weighting factors, and $\mathcal{L}_{\text{base}}$ represents any regularization loss added to improve model generalization.
4 Experiments
4.1 Dataset
We used well-established benchmarks in Sentiment Analysis: SemEval 2014 (14lap, 14res) Pontiki et al. (2014), SemEval 2015 (15res) Pontiki et al. (2015), and SemEval 2016 (16res) Pontiki et al. (2016). These datasets, originally curated for the Semantic Evaluation (SemEval) competitions, are widely used for aspect-based sentiment analysis and extraction tasks Xu et al. (2021).
We employed a data processing framework to construct sentiment-enhanced dependency graphs for ASTE datasets. This framework utilizes SpaCy Honnibal et al. (2020) to extract syntactic dependencies and integrates sentiment scores from SenticNet Cambria et al. (2017) to generate weighted adjacency matrices. These enriched graph representations aim to expand the feature space and reduce biased learning. Along with tokenized sentences and their corresponding aspect-opinion-sentiment triplets, the processed data supports downstream tasks, including model training and evaluation.
Figure 3: F1 Score over epochs for the model trained on 14res, 14lap, 15res, and 16res datasets and tested on the 15res dataset. The results indicate that training on datasets other than 15res, such as 14res can yield an even better performance, which suggests the presence of richer action patterns in other datasets for the model to learn.
We evaluated the conformity of our transition operations in extracting aspects and opinions, with action evaluation results listed in Appendix A.3. Since our approach predicts actions rather than tokens, we assessed how well the model learns action patterns across individual datasets. For instance, as shown in Figure 3, the model did not perform best on the 15res test set when trained on the 15res training set. This suggests that other datasets, such as 14res, may contain richer action patterns for the model to learn. Consequently, we believe that training on a fused dataset provides a more diverse context, enhancing the model’s ability to learn these actions, which we then evaluated on separate test sets.
Figure 4: Growth on different weights for AOPE task on 14lap testing set.
Figure 5: Growth on different weights for ASTE task on 14lap testing set.
4.2 Baselines
We use two baselines: published scores from prior studies and previous SOTA models run on the fused dataset. Regular baselines include classic attention-based methods (e.g., CMLA Wang et al. (2017)), pipeline approaches (Peng-Two-stage Peng et al. (2020)), sequence-to-sequence frameworks (BARTABSA Yan et al. (2021)), sequence-tagging methods (GTS Wu et al. (2020), JET-BERT Xu et al. (2020)), MRC-based models (COM-MRC Zhai et al. (2022), Triple-MRC Zou et al. (2024)), graph-based approaches (EMC-GCN Chen et al. (2022), DGCNAP Li et al. (2023)), and the SOTA MiniConGTS Sun et al. (2024).
In addition to baselines on separate datasets, we selected two state-of-the-art models with publicly available and runnable code: the most recent model, MiniConGTS Sun et al. (2024), and the earlier state-of-the-art for pair extraction, BARTABSA Yan et al. (2021). We used their original configurations and data processing pipelines, modifying only the training data to include all four training sets from 14lap, 14res, 15res, and 16res. The models were then separately evaluated on the four corresponding test sets.
Models 14Res 14Lap 15res 16res
P R F1 P R F1 P R F1 P R F1
CMLA+Wang et al. (2017) $\diamond$ - - 48.95 - - 44.10 - - 44.60 - - 50.00
Peng-two-stage Peng et al. (2020) $\diamond$ - - 56.10 - - 53.85 - - 56.23 - - 60.04
Dual-MRC Mao et al. (2021) $\diamond$ - - 74.93 - - 63.37 - - 64.97 - - 75.71
SpanMlt Zhao et al. (2020) $\bullet$ - - 75.60 - - 68.66 - - 64.68 - - 71.78
BARTABSA Yan et al. (2021) $\diamond$ - - 77.68 - - 66.11 - - 67.98 - - 77.38
MiniConGTS Sun et al. (2024) $\star$ - - 79.60 - - 73.23 - - 73.87 - - 76.29
SOTAs on fused dataset
BARTABSA-fused 75.40 76.76 76.07 72.36 63.40 67.59 88.85 92.77 90.77 87.06 87.32 87.19
MiniConGTS-fused 81.06 78.82 79.92 75.62 71.61 73.56 93.63 92.72 93.17 78.80 81.82 80.28
Ours on fused dataset
Trans-AOPE ($w_{1}$:$w_{2}$ = 1:1) 91.88 80.83 86.00 85.87 74.68 79.89 94.88 88.25 91.45 94.63 89.10 91.80
Trans-AOPE ($w_{1}$:$w_{2}$ = 1:0) 93.80 80.39 86.58 88.60 74.26 80.80 95.07 88.95 91.91 93.10 89.01 91.01
Table 3: Comparison of different models on multiple datasets for AOPE task. Recall and precision values are omitted where they are not provided. Former best scores are underlined, and current best scores are bold. $\diamond$ are retrieved from Yan et al., 2021. $\bullet$ is retrieved from Zhao et al., 2020, and $\star$ are retrieved from Sun et al., 2024
Models 14Res 14Lap 15Res 16Res
P R F1 P R F1 P R F1 P R F1
Peng-Two-stage Peng et al. (2020) $\star$ 43.24 63.66 51.46 38.87 50.38 42.87 48.07 57.51 52.32 46.96 64.24 54.21
BARTABSA Yan et al. (2021) $\diamond$ 65.52 64.99 65.25 61.41 56.19 58.69 59.14 59.38 59.26 66.60 68.68 67.62
GTS Wu et al. (2020) $\star$ 67.76 67.29 67.50 57.82 51.32 54.36 62.59 57.94 60.15 66.08 66.91 67.93
JET-BERT Xu et al. (2020) $\star$ 70.56 55.94 62.40 55.39 43.57 51.04 64.45 51.96 57.53 70.42 58.37 63.83
COM-MRC Zhai et al. (2022) $\star$ 75.46 68.91 72.01 58.15 60.17 61.17 68.35 61.24 64.53 71.55 71.59 71.57
EMC-GCN Chen et al. (2022) $\star$ 71.21 72.39 71.78 61.70 56.26 58.81 61.54 62.47 61.93 65.62 71.30 68.33
BDTF Chen et al. (2022) $\star$ 75.53 73.24 74.35 68.94 55.97 61.74 68.76 63.71 66.12 71.44 73.13 72.27
DGCNAP Li et al. (2023) $\star$ 72.90 68.69 70.72 62.02 53.79 57.57 62.23 60.21 61.19 69.75 69.44 69.58
Triple-MRC Zou et al. (2024) $\star$ - - 72.45 - - 60.72 - - 62.86 - - 68.65
STAGE-3D Liang et al. (2023) $\star$ 78.58 69.58 73.76 71.98 53.86 61.58 73.63 57.90 64.79 76.67 70.12 73.24
MiniConGTS Sun et al. (2024) $\star$ 76.10 75.08 75.59 66.82 60.68 63.61 66.50 63.86 65.15 75.52 74.14 74.83
SOTAs on fused dataset
BARTABSA-fused 71.05 72.33 71.68 63.66 56.01 59.59 88.98 92.77 90.77 83.82 84.07 83.95
MiniConGTS-fused 73.89 72.84 73.36 68.26 57.42 62.37 90.64 89.32 89.98 72.73 73.21 72.97
Ours
Trans-ASTE ($w_{1}$:$w_{2}$ = 1:1) 89.30 77.10 82.76 83.81 69.64 76.07 93.80 86.71 90.11 92.18 86.81 89.42
Trans-ASTE ($w_{1}$:$w_{2}$ = 1:0) 91.70 78.24 84.44 84.63 70.59 76.98 94.31 88.11 91.11 92.78 87.58 90.11
Table 4: Comparison of different models on multiple datasets for ASTE task. Former best scores are underlined, and current best scores are bold. $\star$ are retrieved from Sun et al., 2024, and $\diamond$ is retrieved from Yan et al., 2021.
4.3 Results and Analysis
4.3.1 Study on Augmented Optimization
To assess the impact of weight compositions in Equation 10, we tested $(w_{1},w_{2})\in\{(1:0),(1:1),(1:10),(0:1),(10:1)\}$, where $w_{1}$ and $w_{2}$ are the weights for base and contrastive-based optimization, respectively, using a batch size of 4. Results for AOPE and ASTE are presented in Figures 5 and 5. For AOPE, contrastive-heavy weights ($w_{1}:w_{2}=1:1$ and $w_{1}:w_{2}=1:10$) performed well during early epochs (1–50), demonstrating the early advantages of contrastive optimization. After epoch 50, base ($w_{1}:w_{2}=1:0$) and balanced ($w_{1}:w_{2}=1:1$) optimizations provided greater stability and better final performance. In ASTE, base and balanced optimizations consistently outperformed others, indicating that ASTE primarily benefits from foundational learning, with limited advantages from contrastive optimization.
Similar trends were observed across other datasets, as shown in Appendix A.4, highlighting the broader applicability of these findings. Although contrastive-enhanced optimization does not deliver the best overall performance, it effectively improves efficiency in pair extraction tasks.
4.3.2 Comparison with Baselines
On AOPE
[…]
On ASTE
Results for ASTE are listed in Table 4. For the 14Res dataset, Trans-ASTE achieves the highest F1 score (84.44%), representing an APG of 12.76% over BARTABSA-fused. Similarly, it outperforms MiniConGTS-fused with an APG of 11.08%. In the 14Lap dataset, Trans-ASTE) achieves an APG value of 17.39% in F1 compared to BARTABSA-fused, and 14.61% over MiniConGTS-fused. For the 15Res dataset, Trans-ASTE) maintains a high level of performance, with an F1 score (91.11%) slightly leading both baselines. The overall F1 improvement is 0.34% compared to BARTABSA-fused and 1.13% compared to MiniConGTS-fused. The 16Res dataset also sees consistently high performance from Trans-ASTE), with the highest F1 score (90.11%), representing an APG of 6.16% over BARTABSA-fused.
Overall, combining datasets for all-for-one training does not always result in performance improvements across all datasets for SOTA models. However, when data is fused for our model, it consistently achieves performance gains across datasets, with significant improvements observed in 14Res, 14Lap, and 16Res. That said, the slightly lower performance on the 15Res AOPE task highlights potential areas for further optimization in specific scenarios.
5 Conclusion
In this paper, we present an efficient transition pipeline for aspect-opinion pair extraction with linear time complexity $O(n)$, enhanced by a contrastive learning-based optimization method. This approach eliminates the need to directly identify and extract individual tokens, avoiding token-level bias. It can be trained on datasets with the most comprehensive coverage of the designed actions, resulting in significant performance improvements across various datasets. Specifically, training our model on a well-covered fused dataset enables it to learn robust action combination patterns, leading to superior performance on all datasets. Our model surpasses SOTA models on the same fused dataset, setting new benchmarks for AOPE and ASTE tasks. This approach not only enhances model efficiency but also demonstrates its effectiveness and versatility in improving performance.
Limitation
The primary limitation of the transition-based pipeline lies in its reliance on relatively large and diverse datasets to effectively learn action patterns. As evidenced by the findings in Figure 3, the model achieved better performance on the 15res test set when trained on datasets like 14res and 16res, which appear to provide richer and more diverse action patterns compared to the 15res training set. This observation highlights the inadequacy of relying solely on a single dataset, as it may lack the breadth and complexity needed to capture the full spectrum of action patterns required for robust model generalization. Without access to such diverse training data, the model’s ability to adapt to nuanced patterns or unseen contexts in the test data may be significantly compromised, emphasizing the importance of training on a fused or broader dataset to enhance its overall effectiveness.
Appendix A Appendix
A.1 Alternative Solution with State Changes
Since stack can hold more than two tokens, "is" can be moved to stack ($\beta_{3}\rightarrow\sigma_{3}$) before "Gourmet" and "food" merge ($[\sigma_{1},\sigma_{2},\sigma_{3}]\rightarrow[\sigma_{1\&2},\sigma_{3}]$). The detailed process is present in Table 5.
Phrase Action Stack ($\sigma$) Buffer ($\beta$) Aspect Opinion Pair
[] [$\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$]
1 $SF$ [$\sigma_{1}$] [$\beta_{2}$, $\beta_{3}$, $\beta_{4}$]
2 $SF$ [$\sigma_{1}$, $\sigma_{2}$] [$\beta_{3}$, $\beta_{4}$]
3 $SF$ [$\sigma_{1}$, $\sigma_{2}$, $\beta_{3}$] [$\beta_{4}$]
4 $M$ [$\sigma_{1\&2}$, $\beta_{3}$] [$\beta_{4}$]
5 $R_{n}$ [$\sigma_{1\&2}$] [$\beta_{4}$]
6 $SF$ [$\sigma_{1\&2}$, $\sigma_{4}$] []
7 $RR$ [$\sigma_{1\&2}$, $\sigma_{4}$] [] [$\sigma_{1\&2}$] [$\sigma_{4}$] ($\sigma_{1\&2}\rightarrow\sigma_{4}$)
9 $ST$ [] [] [$\sigma_{1\&2}$] [$\sigma_{4}$] ($\sigma_{1\&2}\rightarrow\sigma_{4}$)
Table 5: State changes for "Gourmet food is delicious" using symbolic representation. Here, $\sigma_{1}$ corresponds to "Gourmet", $\sigma_{2}$ to "food", $\sigma_{3}$ to "is", and $\sigma_{4}$ to "delicious". Similarly, $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, and $\beta_{4}$ correspond to tokens in the buffer in sequence.
A.2 Algorithm for Transition Actions
In Algorithm 1, while the core of the algorithm focuses on the action operations, the configuration also plays a vital role in its functionality.
The algorithm begins with a buffer, a list of input tokens representing the text to be parsed. These tokens are consumed during the parsing process and form the basis for token pair relationships. A critical component of the algorithm is the pair2sent mapping, which associates token pairs with their corresponding sentiment labels, allowing the algorithm to enrich the output with sentiment information, even for token pairs not explicitly provided in the target data (defaulting to a neutral label). Then, each parsing action (e.g., Shift, Left Relation, Right Relation) is mapped to a unique numeric ID. This mapping enables a compact representation of the sequence of parsing actions, which is useful for downstream tasks like training or evaluation. Throughout the parsing process, the algorithm records intermediate states (stack, buffer, actions, and sentiment labels) to maintain a complete history of its operations. Simultaneously, valid token pairs are stored in pairs, capturing the structured relationships in the text.
Algorithm 1 Argument-Opinion Pair Parsing Algorithm
1:buffer (list of input tokens), pair2sent (mapping of token pairs to sentiment labels), action2id (mapping of parsing actions to numeric IDs)
2:states (list of recorded parsing states), pairs (list of valid token pairs)
3:Initialize empty stack, actions, states, and pairs. $\triangleright$ Main processing loop: parse tokens from the buffer
4:while buffer is not empty do
5:     if len(stack) < 2 then
6:         stack.append(buffer.pop(0)) $\triangleright$ Apply $SH$: Shift
7:         Append $SF$ to actions.
8:         continue
9:     end if $\triangleright$ Form a pair from the top two elements in stack
10:     stackPair $\leftarrow$ (stack[-2], stack[-1])
11:     currentAction $\leftarrow$ getAction(stackPair, targetPairs)
12:     Append currentAction to actions. $\triangleright$ Handle relation formation actions ($LR$ and $RR$)
13:     if currentAction == $LR$ then
14:         Add stackPair to pairs.
15:         sentimentLabel $\leftarrow$ pair2sent.get(stackPair, 3)
16:         Apply stack[-2] and stack[-1] form a left-relation pair.
17:     else if currentAction == $RR$ then
18:         Add stackPair to pairs.
19:         sentimentLabel $\leftarrow$ pair2sent.get(stackPair, 3)
20:         Apply stack[-1] and stack[-2] form a right-relation pair.
21:     else
22:         sentimentLabel $\leftarrow 3$triangleright Default sentiment label.
23:     end if $\triangleright$ Record the current parsing state
24:     Record (stack, buffer, actions, sentimentLabel) in states. $\triangleright$ Modify the stack based on currentAction
25:     if currentAction == $L_{n}$ then
26:         Remove stack[-2] $\triangleright$ Apply $L_{n}$: Left Constituent Removal
27:     else if currentAction == $R_{n}$ then
28:         Remove stack[-1] $\triangleright$ Apply $R_{n}$: Right Constituent Removal
29:     else if currentAction == $M$ then
30:         Merge stack[-2] and stack[-1] $\triangleright$ Apply $M$: Merge
31:     end if
32:end while $\triangleright$ Finalize by reducing the remaining elements in the stack
33:while len(stack) > 1 do
34:     Remove the top element from stack $\triangleright$ Apply $ST$: Stop
35:end while $\triangleright$ Convert actions to numeric IDs for output
36:Map each action in states to its ID using action2id. return states, pairs.
A.3 Action Coverage on Different Datasets
14lap 14res 15res 16res fused
train (Precision) 99.99 99.99 99.99 99.99 99.99
train (Recall) 85.51 86.63 92.98 93.04 88.73
train (F1) 92.19 92.84 96.36 96.40 94.03
dev (Precision) 99.99 99.99 99.99 99.99 99.99
dev (Recall) 85.97 86.89 94.92 93.04 89.88
dev (F1) 92.46 92.99 97.39 96.40 94.67
test (Precision) 99.99 99.99 99.99 99.99 99.99
test (Recall) 86.66 87.54 91.47 92.42 89.41
test (F1) 92.85 93.36 95.54 96.06 94.41
total (Precision) 99.99 99.99 99.99 99.99 99.99
total (Recall) 86.02 87.03 92.56 92.79 89.00
total (F1) 92.48 93.06 96.14 96.26 94.18
Table 6: Transposed Transitional Action Coverage for 14Lap, 14Res, 15Res, 16Res, and fused Datasets.
The coverage of actions, as shown in Table 6, reveals some differences in the recall values across data sets. The 14Lap and 14Res datasets exhibit lower recall compared to the 15Res and 16Res datasets. This trend is observed across training, development, and testing phases, with recall values for 14 datasets remaining around 85-87%, while 15Res and 16Res achieve significantly higher recalls, exceeding 91% in most cases. The fused dataset strikes a balance with intermediate recall values, reflecting the integration of characteristics from all datasets. This highlights a relative weakness in capturing relevant actions in 14 datasets compared to 15 and 16 datasets.
A.4 Growth with Different Optimization Weight
Figures 7 and 7, 9 and 9, and 11 and 11 depict the performance growth with varying weights applied to the 14res, 15res, and 16res datasets, respectively. These figures highlight the influence of weight adjustments on model accuracy and efficiency in ASTE tasks. By examining the trends across these datasets, we demonstrate the early benefits of augmented optimization in enhancing performance.
Figure 6: Growth on different weights for AOPE task on 14res testing set.
Figure 7: Growth on different weights for ASTE task on 14res testing set.
Figure 8: Growth on different weights for AOPE task on 15res testing set.
Figure 9: Growth on different weights for ASTE task on 15res testing set.
Figure 10: Growth on different weights for AOPE task on 16res testing set.
Figure 11: Growth on different weights for ASTE task on 16res testing set.
A.5 Growth Comparison for AOPE and ASTE
The growth comparison for AOPE and ASTE is illustrated in Figures 13, 13, 15, and 15, corresponding to the 14res, 14lap, 15res, and 16res datasets, respectively. While ASTE generally follows the trends observed in AOPE, the fluctuations in ASTE are more extreme. These figures emphasize the sensitivity of ASTE to variations in model weights or dataset conditions, which results in sharper peaks and deeper troughs compared to the relatively smoother patterns seen in AOPE. This highlights the unique challenges posed by the ASTE task across different datasets.
Figure 12: Growth Comparison for AOPE and ASTE on 14res Testing set.
Figure 13: Growth Comparison for AOPE and ASTE on 14lap Testing set.
Figure 14: Growth Comparison for AOPE and ASTE on 15res Testing set.
Figure 15: Growth Comparison for AOPE and ASTE on 16res Testing set.